OpenStudy (yrivers36):

Factor the expression if possible. (x+1)^2-25 with explanation please.

OpenStudy (jamesj):

So expand this out first and what do you get?

OpenStudy (yrivers36):

not quite sure what you are asking

OpenStudy (yrivers36):

(x+1)(x+1)

OpenStudy (jamesj):

(x+1)^2-25 = x^2 + 2x + 1 - 25 = x^2 + 2x - 24

OpenStudy (jamesj):

Now factor that by finding the roots.

OpenStudy (anonymous):

why not just factor it as a difference of two squares; then simplify

OpenStudy (anonymous):

\[a^2−b^2=(a−b)(a+b)\] where a=x+1 and b=5

OpenStudy (jamesj):

Sure, that too.

OpenStudy (anonymous):

\[(x+1)^2-5^2=(x+1-5)(x+1+5)=...\]

OpenStudy (yrivers36):

Ok I like it the way JamesJ did it...Mandoline it seems a little to complicated for me

OpenStudy (jamesj):

Mandolino's way is actually more elegant, but a bit more confusing for beginners.

OpenStudy (yrivers36):

oh ok